# Detection of Correlated Random Vectors

### v1

@@ 4 Lower Bound via Polynomial Decomposition
This section is devoted to the proofs of Theorems 12. The proof strategy for our lower bound begins with a series of standard arguments frequently used in proofs of impossibility of detection in hypothesis testing problems involving a planted combinatorial structure. We start by recalling the well-known characterization of the optimal Bayes risk, using the total-variation distance, i.e.,
@@
[…]
@@
The above equation shows that if $\mathbb{E}_{{\cal H}_{0}}[{\cal L}^{2}]$ decays, then weak and strong detection are impossible.
At this point, we deviate from the typical straight-forward analysis of the second moment of the likelihood ratio as performed in [EH23, PH23, KN22b], which involves a non-trivial analysis of the cycles structure of random permutations. Not only that the straight-forward analysis is highly complicated, as we show, it also results in a non-tight lower bound. We suggest a simple yet powerful alternative technique for bounding the second moment of the likelihood ratio, using a decomposition of ${\cal L}(\mathsf{X},\mathsf{Y})$ to orthogonal components w.r.t. the orthonormal base of Hermite polynomials.
Consider the space $L^{2}({\cal H}_{0})$, of all real valued random variables $\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ with finite second moment (w.r.t. the measure ${\mathbb{P}}_{{\cal H}_{0}}$), which is a Hilbert space when equipped with the inner product given by
$\left\langle\phi,\psi\right\rangle_{{\cal H}_{0}}\triangleq\mathbb{E}_{{\cal H}_{0}}\left[\psi(\mathsf{X},\mathsf{Y})\cdot\phi(\mathsf{X},\mathsf{Y})\right].$ (26)
We recall the univariate Hermite polynomials, a sequence polynomials $(h_{k}(x))_{k\geq 0}$, with $\mathsf{deg}(h_{k})=k$, defined by the equation
$h_{k}(x)=(-1)^{k}e^{x^{2}/2}\frac{d^{k}}{dx^{k}}e^{-x^{2}/2}.$ (27)
It is well-known (see, for example [MOS66][Chapter 5.6]) that Hermite polynomials are orthonormal w.r.t. the standard Gaussian measure. Namely,
The multivariate Hermite polynomials in $n$ variables are indexed by $\theta\in\mathbb{N}^{n}$, and are merely products of the univariate Hermite polynomials: $H_{\theta}(x)=\prod_{i=1}^{n}h_{\theta_{i}}(x_{i})$. We note that the degree of $H_{\theta}$ is exactly $|\theta|\triangleq\sum_{i}\theta_{i}$. The multivariate Hermite polynomials $(H_{\theta})_{\theta\in\mathbb{N}^{n}}$ are orthonormal w.r.t. an independent product of standard Gaussian measures:
$\mathbb{E}_{Y\sim N(\mathbf{0},\mathbf{I})}\left[H_{\alpha}(Y)H_{\gamma}(Y)\right]=\delta[\alpha-\gamma]$ (29)
The key idea of our technique is to decompose the likelihood ratio to its orthogonal components using the basis composed from the (infinite) family of Hermite polynomial, and then to use Parseval’s equality. Formally, since the Hermitian polynomials compose a complete orthonormal system in $L^{2}({\cal H}_{0})$ (see for example [UR19, Proposition 1.10]), ${\cal L}(\mathsf{X},\mathsf{Y})$ is decomposed as
@@
which converges in norm to ${\cal L}$ and by Parseval’s equality
@@
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}.$ (34)
In the following Lemma, we calculate the projection coefficients in the above sum.
@@
For any $\alpha,\beta\in\mathbb{N}^{n}$,
@@
We begin with a simple observation - by the definition of the Radon-Nikodym derivative we have
@@
where the last equality follows from the independence of the pairs $(\mathsf{X}_{i},\mathsf{Y}_{\sigma(i)})_{i=1}^{n}$ given the hidden permutation $\sigma$. We recall that given the permutation $\sigma$, the $(\mathsf{X}_{i},\mathsf{Y}_{\sigma(i)})$ has a two dimensional normal distribution ${\cal N}(\mathbf{0},\Sigma_{\rho})$. Let $X$ and $Z$ be two independent standard normal random variables. A simple calculation shows that the pair $(X,\rho X+\sqrt{1-\rho^{2}}Z)$ is also distributed as ${\cal N}(\mathbf{0},\Sigma_{\rho})$, and therefore for all $i$,
@@
where the last equality is by the law of total expectation. We conditioned on $X$, the random variable $\rho X+\sqrt{1-\rho^{2}}Z$ is distributed as ${\cal N}(\rho X,1-\rho^{2})$. Thus, our goal now is to evaluate the expectation $\mathbb{E}_{Y\sim{\cal N}(\rho x,1-\rho^{2})}[h_{k}(Y)]$, for any $k\in\mathbb{N}$ and $|\rho|<1$.
Let $\bar{h}_{k}$ denote the $k$-th order physicist’s Hermite polynomials. It is known that [Bac]
@@
for $|\rho|<1$, $k\geq 0$, and $y\in\mathbb{R}$. Also, it is known that [MOS66, Chapter 5.6] $h_{k}(x)=2^{-\frac{k}{2}}\bar{h}_{k}\left(\frac{x}{\sqrt{2}}\right)$. Thus,
@@
$=\frac{2^{-\frac{k}{2}}}{\sqrt{\pi(1-\rho^{2})}}\sqrt{\pi}\rho^{k}\sqrt{1-\rho^{2}}\bar{H}_{k}(x/\sqrt{2})$ (48)
@@
Combining (43) and (50) now have
$\mathbb{E}_{{\cal H}_{1}|\sigma}\left[h_{\alpha_{i}}(\mathsf{X}_{i})\cdot h_{\beta_{\sigma(i)}}(\mathsf{Y}_{\sigma(i)})\right]$ $=\mathbb{E}\left[h_{\alpha_{i}}(X)\cdot\rho^{\beta_{\sigma(i)}}\cdot h_{\beta_{\sigma(i)}}(X)\right]$ (51)
$=\rho^{\beta_{\sigma(i)}}\cdot\delta[\alpha_{i}-\beta_{\sigma(i)}].$ (52)
We now conclude,
@@
$=\rho^{|\alpha|}{\mathbb{P}}[\sigma(\beta)=\alpha],$ (56)
where the last equality follows since if ${\mathbb{P}}[\sigma(\beta)=\alpha]$ is non-zero, in particular $|\alpha|=|\beta|$. ∎
We now turn to analyse ${\mathbb{P}}[\sigma(\beta)=\alpha]$ for arbitrary vectors $\alpha,\beta\in\mathbb{N}^{n}$. To that end, we consider the integer distribution function of a vector $\alpha\in\mathbb{N}^{n}$, $p_{\alpha}:\mathbb{N}\to\mathbb{N}$, where $p_{\alpha}(\ell)$ is defined to be the number of coordinates of $\alpha$ that contain $\ell$. We also consider the equivalence relation over $\mathbb{N}^{n}$ defined by setting $\alpha\equiv\beta$ if and only if there is $\sigma\in{\mathbb{S}}_{n}$ such that $\sigma(\beta)=\alpha$. We easily observe that $\alpha\equiv\beta$ if and only if the have the same distribution functions, namely, $p_{\alpha}=p_{\beta}$. Let $[\alpha]$ denote the equivalence class of $\alpha$ w.r.t. this equivalence relation. For $\sigma\sim\mathsf{Unif}({\mathbb{S}}_{n})$, symmetry implies that $\sigma(\alpha)\sim\mathsf{Unif}([\alpha])$. We therefore have
@@
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}$ (60)
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n}\\ |\alpha|+|\beta|=m\end{subarray}}\frac{\rho^{2m}}{|[\alpha]|^{2}}\mathbbm{1}_{\alpha\equiv\beta}$ (63)
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n}\\ |\alpha|=|\beta|=m\end{subarray}}\frac{\rho^{2m}}{|[\alpha]|^{2}}\mathbbm{1}_{\alpha\equiv\beta}$ (66)
@@
=$\overset{(a)}{=}\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}[\alpha]\\ \left|\alpha\right|=m\end{subarray}}\sum_{\beta\in[\alpha]}\frac{\rho^{2|\beta|}}{|[\beta]|}$ (80)
=$\overset{(b)}{=}\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}[\alpha]\\ \left|\alpha\right|=m\end{subarray}}\rho^{2m}$ (83)
@@
where $(a)$ follows since the set of equivalence classes is a partition of $\mathbb{N}^{n}$, and $(b)$ follows since $[\beta]=[\alpha]$ for $\alpha\equiv\beta$.
For a fixed $m\in\mathbb{N}$, consider the set of all equivalence classes $([\alpha])_{\alpha\in\mathbb{N}^{n}}$ with $\left|\alpha\right|=m$ (which is well defined as the $L^{1}$ norm is preserved under equivalence relation $\equiv$). As we mentioned, $\alpha\equiv\beta$ if and only if $p_{\alpha}=p_{\beta}$. We also note that for any $\alpha$, the corresponding distribution function $p_{\alpha}$ the sum $\sum_{k\in\mathbb{N}}p_{\alpha}(k)\leq n$ is the number of non-zero elements of $\alpha$, and the sum $\sum_{k\in\mathbb{N}}k\cdot p_{\alpha}(k)=|\alpha|$ is the degree of the polynomial $H_{\alpha}$. Thus, equivalence classes of the form $([\alpha])_{\alpha\in\mathbb{N}^{n}}$, $|\alpha|=m$, are in a one-to-one correspondence with integer distribution functions $p:\mathbb{N}\to\mathbb{N}$ which satisfy
@@
Such integer distributions represents integer partitions, which are defined as follows:
@@
Indeed, any distribution function which satisfy (85) correspond to a unique partition $\mathcal{p}$ of $m$ to at most $n$ elements given by the multiset in which any integer $k\in\mathbb{N}$ has exactly $p(k)$ copies. Hence, we conclude that
|$|\left\{[\alpha]:|\alpha|=m\right\}|$ =|$=\left|\left\{p:\mathbb{N}\to\mathbb{N}~{}\big{|}~{}\sum_{k\in\mathbb{N}}k\cdot p(k)=m,\sum_{k\in\mathbb{N}}p(k)=n\right\}\right|$ (86)
$=\left|\mathsf{Par}(m,\leq n)\right|.$ (87)
We are ready to prove our lower bounds. Combining the above equality (87) with (84) we bound the second moment of the likelihood ratio
@@
$=\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{n})\right|\rho^{2m},$ (89)
Since $\left|\mathsf{Par}(m,\leq n)\right|$ is monotone non-decreasing with $n$ we have
$\mathbb{E}_{{\cal H}_{0}}[{\cal H}(\mathsf{X},\mathsf{Y})]\leq\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|\rho^{2m},$ (90)
where $|\mathsf{Par}(m,\leq_{\infty})|$ is just the number of integer partitions of the number $m$ (without limitations on the number of parts).
By the well-known Hardy-Ramanujan Formula for the number of integer partitions, there exists some universal constant $c\in\mathbb{R}$ such that for all $m\in\mathbb{N}$
$|\mathsf{Par}(m,\leq_{\infty})|\leq c\cdot\frac{1}{4\sqrt{3}m}\exp\left(\pi\sqrt{\frac{2m}{3}}\right),$ (91)

### v3

@@ 4 Lower Bound via Polynomial Decomposition
This section is devoted to the proofs of Theorems 12. A crucial part of our proof involves the expansion of the likelihood function into its orthogonal Hermite polynomials components, which compose an orthonormal basis of $L^{2}$ w.r.t. the Gaussian measure. We shall therefore begin with a brief introduction of Hermite polynomials in Hilbert spaces. For the benefit of readability, throughout this entire section, with some abuse of notation, we use lower-case Greek letters (e.g., $\alpha$), for integer vectors of finite lengths.
4.1 Hermite polynomials and Hilbert spaces
Let $\mathsf{Y}\sim{\cal N}(0,\mathbf{I})$ be an $n$-dimensional random variable distributed according to, $\mu$, an $n$-dimensional Gaussian distribution. Consider the space $L^{2}(\mu)$, of real-valued random variables $\mathbb{R}^{n}\to\mathbb{R}$ with finite second moment (w.r.t. the measure to the measure $\mu$), which is a Hilbert space when equipped with the inner product,
$\left\langle\phi,\psi\right\rangle_{\mu}\triangleq\mathbb{E}_{\mu}\left[\phi(\mathsf{Y})\cdot\psi(\mathsf{Y})\right],$ (16)
where $\phi$ and $\psi$ are two measurable functions from $\mathbb{R}^{n}$ to $\mathbb{R}$. Recall the univariate Hermite polynomials, which are a sequence of polynomials $(h_{k}(x))_{k\geq 0}$, with $\mathsf{deg}(h_{k})=k$, defined by the equation,
$h_{k}(x)\triangleq(-1)^{k}e^{x^{2}/2}\frac{d^{k}}{dx^{k}}e^{-x^{2}/2}.$ (17)
It is well-known (see, for example [25][Chapter 5.6]) that Hermite polynomials are orthonormal w.r.t. the standard one-dimensional Gaussian measure. Namely,
The multivariate Hermite polynomials in $n$ variables are indexed by $\theta\in\mathbb{N}^{n}$, and are merely products of the univariate Hermite polynomials, i.e., $H_{\theta}(x)=\prod_{i=1}^{n}h_{\theta_{i}}(x_{i})$. We note that the degree of $H_{\theta}$ is exactly $|\theta|\triangleq\sum_{i}\theta_{i}$, which is exactly the $L_{1}$ norm of the vector $\theta$. Furthermore, the multivariate Hermite polynomials $(H_{\theta})_{\theta\in\mathbb{N}^{n}}$ are orthonormal w.r.t. an independent product of standard Gaussian measures, namely,
$\mathbb{E}_{\mathsf{Y}\sim N(\mathbf{0},\mathbf{I})}\left[H_{\alpha}(\mathsf{Y})H_{\gamma}(\mathsf{Y})\right]$ $=\mathbb{E}_{\mathsf{Y}\sim N(\mathbf{0},\mathbf{I})}\left[\prod_{i=1}^{n}h_{\alpha_{i}}(Y_{i})h_{\gamma_{i}}(Y_{i})\right]$ (19)
=$\overset{(a)}{=}\prod_{i=1}^{n}\mathbb{E}_{Y_{i}\sim N(\mathbf{0},1)}\left[h_{\alpha_{i}}(Y_{i})h_{\gamma_{i}}(Y_{i})\right]$ (20)
=$\overset{(b)}{=}\prod_{i=1}^{n}\delta[\alpha_{i}-\gamma_{i}]=\delta[\alpha-\gamma],$ (21)
where $(a)$ follows from the independence of the random variables $\left\{Y_{i}\right\}_{i}$, and $(b)$ follows from (16).
It is well-known that the Hermitian polynomials form a complete orthonormal basis in the Hilbert space $L^{2}(\mu)$ w.r.t. the Gaussian measure (see, e.g., [26, Proposition 1.10]), and therefore any random variable of the form $\psi(\mathsf{Y})$ can be expanded as,
$\psi(\mathsf{Y})=\sum_{\theta\in\mathbb{N}^{n}}\left\langle H_{\theta}(\mathsf{Y}),\psi(\mathsf{Y})\right\rangle_{\mu}H_{\theta}(\mathsf{Y}),$ (22)
where the equality is in the sense that the sum at the r.h.f.s of (22) converges in norm to $\psi(\mathsf{Y})$. Finally, Parseval’s identity implies that,
$\mathbb{E}_{\mu}\left[\psi(\mathsf{Y})^{2}\right]$ $=\left\|\psi(\mathsf{Y})\right\|^{2}_{\mu}$ (23)
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\theta\in\mathbb{N}^{n}\\ |\theta|=m\end{subarray}}\left|\left\langle H_{\theta}(\mathsf{Y}),\psi(\mathsf{Y})\right\rangle_{\mu}\right|^{2}.$ (24)
4.2 Proof outline
Before delving into the proofs, let us provide a brief outline of the main steps. Roughly speaking, we derive our lower bounds using a non-standard analysis of the the second moment of likelihood function which involves the decomposition of the likelihood function into its orthogonal components, which we presented in the previous subsection. The main steps are:
  1. 1.
    Bounding the risk using the likelihood’s second moment: Using a series of fairly standard arguments, we bound the optimal risk function as
  2. 2.
    Hermite polynomial decomposition: Using Parseval’s identity we present the likelihood’s second moment (with respect to ${\cal H}_{0}$) using the orthogonal projection coefficients,
    $\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]=\sum_{(\alpha,\beta)\in\mathbb{N}^{2n}}\left|\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}\right|^{2},$ (26)
    where $H_{\alpha,\beta}$ is the Hermite polynomial with $2n$ variables which corresponds to the pair $(\alpha,\beta)\in\mathbb{N}^{2n}$.
  3. 3.
    Computing the orthogonal projection: In Lemma 1, we show that the coefficients of the orthogonal projection coefficients can be presented as,
    $\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}=\mathbb{E}_{\sigma}\left[\mathbb{E}_{{\cal H}_{1}|\sigma}\left[H_{\alpha,\beta}(\mathsf{X},\mathsf{Y})\right]\right].$ (27)
    Then, we carefully analyse the r.h.s. of the above equation using some well-known properties of Hermite polynomials and give an expression which depends only on $\alpha$ and $\beta$.
  4. 4.
    Integer partitions: We plug in the result in the previous step into (26), and show that the likelihood’s second moment is uniformly bounded by the generating function of integer partitions (see, Definition 2). Finally, we combine this bound and (25) in order to get a threshold for the impossibility of detection.
4.3 The proof of the lower bound
@@
The proof strategy for our lower bound begins with a series of standard arguments frequently used in proofs of impossibility of detection in hypothesis testing problems involving a planted combinatorial structure. We start by recalling the well-known characterization of the optimal Bayes risk, using the total-variation distance [27, Theorem 2.2],
@@
$d_{\mathsf{TV}}({\mathbb{P}}_{{\cal H}_{0}},{\mathbb{P}}_{{\cal H}_{1}})=\frac{1}{2}\intop\left|{\cal L}(\mathsf{X},\mathsf{Y})-1\right|\mathrm{d}{\mathbb{P}}_{{\cal H}_{0}},$ (29)
@@
$\mathsf{R}^{\star}$ $\geq 1-\frac{1}{2}\sqrt{\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]-1}$ (31)
$=1-\frac{1}{2}\sqrt{\chi^{2}({\mathbb{P}}_{{\cal H}_{0}},{\mathbb{P}}_{{\cal H}_{1}})},$ (32)
where $\chi^{2}$ is the $\chi^{2}$-divergence/distance [27]. The above equation shows that if $\mathbb{E}_{{\cal H}_{0}}[{\cal L}^{2}]$ approaches to $1$, then weak and strong detection are impossible.
At this point, we deviate from the typical straight-forward analysis of the second moment of the likelihood ratio as in [6, 7, 4], which involves a non-trivial analysis of the cycles structure of random permutations. Not only that this analysis is highly complicated, as we show, it also results in a loose lower bound. Instead, we suggest a simple, but powerful alternative technique for bounding the second moment of the likelihood ratio, using a decomposition/expansion of the likelihood function ${\cal L}(\mathsf{X},\mathsf{Y})$ w.r.t. the orthonormal basis of Hermite polynomials.
The key idea of our technique is to decompose the likelihood ratio into its orthogonal components using the basis formed by the (infinite) family of Hermite polynomials, and then use Parseval’s identity. As mentioned in Section 4.1, the Hermitian polynomials form a complete orthonormal system in $L^{2}({\cal H}_{0})$, and by (24), ${\cal L}(\mathsf{X},\mathsf{Y})$ can be expanded as,
@@
where $H_{\alpha,\beta}$ denotes the Hermite polynomial with $2n$ variables defined by $(\alpha,\beta)\in\mathbb{R}^{2n}$. The above sum converges in norm to ${\cal L}$, and by Parseval’s identity,
@@
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in\mathbb{N}^{2n}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}.$ (35)
In the following lemma, we find an expression for the projection coefficients in (35).
@@
For any $(\alpha,\beta)\in\mathbb{N}^{2n}$,
@@
By the definition of the Radon-Nikodym derivative we have,
@@
where the last equality follows from the independence of the pairs $(\mathsf{X}_{i},\mathsf{Y}_{\sigma(i)})_{i=1}^{n}$ conditioned on the hidden permutation $\sigma$. Given $\sigma$, the pair of random variables $(\mathsf{X}_{i},\mathsf{Y}_{\sigma(i)})$ is distributed as the two-dimensional Gaussian distribution ${\cal N}(\mathbf{0},\Sigma_{\rho})$. Let $X$ and $Z$ be two independent standard normal random variables. A simple calculation reveals that the pair $(X,\rho X+\sqrt{1-\rho^{2}}Z)$ is also distributed as ${\cal N}(\mathbf{0},\Sigma_{\rho})$, and therefore for all $i$,
@@
where the last equality follows from the law of total expectation. When conditioned on $X$, the random variable $\rho X+\sqrt{1-\rho^{2}}Z$ satisfies
$(\rho X+\sqrt{1-\rho^{2}}Z)~{}\big{|}~{}X\sim{\cal N}(\rho X,1-\rho^{2}).$ (45)
Thus, our goal is now to find the expectation $\mathbb{E}_{Y\sim{\cal N}(\rho x,1-\rho^{2})}[h_{k}(Y)]$, for any $k\in\mathbb{N}$ and $|\rho|<1$. Let $\bar{h}_{k}$ denote the $k$-th order physicist’s Hermite polynomials. It is known that [28, equation (3)],
@@
for $|\rho|<1$, $k\geq 0$, and $y\in\mathbb{R}$. Also, it is known that [25, Chapter 5.6]
$h_{k}(x)=2^{-\frac{k}{2}}\bar{h}_{k}\left(\frac{x}{\sqrt{2}}\right).$ (47)
Thus,
@@
=$\overset{(a)}{=}\frac{2^{-\frac{k}{2}}}{\sqrt{\pi(1-\rho^{2})}}\sqrt{\pi}\rho^{k}\sqrt{1-\rho^{2}}\bar{h}_{k}(x/\sqrt{2})$ (51)
@@
where $(a)$ follows by using (46) with $\tilde{x}=x/\sqrt{2}$. Combining (44), (45), and (53) we obtain,
$\mathbb{E}_{{\cal H}_{1}|\sigma}\left[h_{\alpha_{i}}(\mathsf{X}_{i})\cdot h_{\beta_{\sigma(i)}}(\mathsf{Y}_{\sigma(i)})\right]$ $=\mathbb{E}\left[h_{\alpha_{i}}(X)\cdot\mathbb{E}\left[h_{\beta_{\sigma(i)}}(\rho X+\sqrt{1-\rho^{2}}Z)~{}\big{|}~{}X\right]\right]$ (54)
$=\mathbb{E}\left[h_{\alpha_{i}}(X)\cdot\mathbb{E}_{Y\sim N(\rho X,1-\rho^{2})}\left[h_{\beta_{\sigma(i)}}(Y)\right]\right]$ (55)
$=\mathbb{E}\left[h_{\alpha_{i}}(X)\cdot\rho^{\beta_{\sigma(i)}}\cdot h_{\beta_{\sigma(i)}}(X)\right]$ (56)
$=\rho^{\beta_{\sigma(i)}}\cdot\delta[\alpha_{i}-\beta_{\sigma(i)}],$ (57)
where the last equality follows from the orthogonality property of the Hermite polynomials. We now conclude,
@@
$=\rho^{|\beta|}{\mathbb{P}}[\sigma(\beta)=\alpha].$ (61)
Note that the if ${\mathbb{P}}[\sigma(\beta)=\alpha]\neq 0$ then it must be that $|\alpha|=|\beta|$ (since the $L_{1}$ norm is invariant under reordering of the coordinates). Thus, we have
$\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}\rho^{|\beta|}={\mathbb{P}}[\sigma(\beta)=\alpha]=\rho^{|\alpha|}{\mathbb{P}}[\sigma(\beta)=\alpha].$ (62)
We now turn to analyse ${\mathbb{P}}[\sigma(\beta)=\alpha]$ for arbitrary vectors $\alpha,\beta\in\mathbb{N}^{n}$. To that end, we consider the integer distribution function of a vector $\alpha\in\mathbb{N}^{n}$, $p_{\alpha}:\mathbb{N}\to\mathbb{N}$, where $p_{\alpha}(\ell)$ is defined as the number of coordinates of $\alpha$ that contain $\ell$. We also consider the equivalence relation over $\mathbb{N}^{n}$ defined by setting $\alpha\equiv\beta$ if and only if there is $\sigma\in{\mathbb{S}}_{n}$ such that $\sigma(\beta)=\alpha$. We observe that $\alpha\equiv\beta$ if and only if they have the same distribution functions, namely, $p_{\alpha}=p_{\beta}$. Let $[\alpha]$ denote the equivalence class of $\alpha$ w.r.t. this equivalence relation. [^1] For mathematical correctness, we can think of []delimited-[][$\cdot$][ ⋅ ] as a map between integer vectors and equivalence classes, which maps any $\alpha$ to the unique equivalence class for which it belongs. For $\sigma\sim\mathsf{Unif}({\mathbb{S}}_{n})$, symmetry implies that $\sigma(\alpha)\sim\mathsf{Unif}([\alpha])$. We therefore have,
@@
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in\mathbb{N}^{2n}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}$ (64)
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in\mathbb{N}^{2n}\\ |\alpha|+|\beta|=m\end{subarray}}\frac{\rho^{m}}{|[\alpha]|^{2}}\mathbbm{1}_{\alpha\equiv\beta}$ (65)
=$\overset{(a)}{=}\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in\mathbb{N}^{2n}\\ |\alpha|=|\beta|=m\end{subarray}}\frac{\rho^{2m}}{|[\alpha]|^{2}}\mathbbm{1}_{\alpha\equiv\beta}$ (66)
@@
=$\overset{(b)}{=}\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}[\alpha]\\ \left|\alpha\right|=m\end{subarray}}\sum_{\beta\in[\alpha]}\frac{\rho^{2|\beta|}}{|[\beta]|}$ (70)
=$\overset{(c)}{=}\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}[\alpha]\\ \left|\alpha\right|=m\end{subarray}}\rho^{2m}$ (71)
@@
where:
  • $(a)$
    follows since the $L_{1}$ norm is invariant under the equivalence relation $\equiv$ (since reordering of the coordinate do not affect the $L_{1}$ norm) and therefore the summand in (65) may be nonzero only if $|\alpha|=|\beta|$. Thus, the summation includes only pairs $(\alpha,\beta)$ such that $|\alpha|=|\beta|=m$ for some $m\in\mathbb{N}$.
  • $(b)$
    follows since the set of equivalence classes is a partition of $\mathbb{N}^{n}$,
  • $(c)$
    follows since $[\beta]=[\alpha]$ for $\alpha\equiv\beta$.
For a fixed $m\in\mathbb{N}$, consider the set of all equivalence classes $([\alpha])_{\alpha\in\mathbb{N}^{n}}$ with $\left|\alpha\right|=m$ (which is well defined as the $L^{1}$ norm is preserved under equivalence relation $\equiv$). As we mentioned, $\alpha\equiv\beta$ if and only if $p_{\alpha}=p_{\beta}$. We also note that for any $\alpha$, with corresponding distribution function $p_{\alpha}$, the sum $\sum_{k\in\mathbb{N}}p_{\alpha}(k)\leq n$ is the number of non-zero elements of $\alpha$, and the sum $\sum_{k\in\mathbb{N}}k\cdot p_{\alpha}(k)=|\alpha|$ is the degree of the polynomial $H_{\alpha}$. Thus, equivalence classes of the form $([\alpha])_{\alpha\in\mathbb{N}^{n}}$, $|\alpha|=m$, are in a one-to-one correspondence with integer distribution functions $p:\mathbb{N}\to\mathbb{N}$ which satisfy
@@
Such integer distributions represent integer partitions, which are defined as follows.
@@
Indeed, any distribution function which satisfy the conditions in (73) correspond to a unique partition $\mathcal{p}$ of $m$ to at most $n$ elements given by the multiset in which any integer $k\in\mathbb{N}$ has exactly $p(k)$ copies. Hence, we conclude that,
|$|\left\{[\alpha]:|\alpha|=m\right\}|$ =|$=\left|\left\{p:\mathbb{N}\to\mathbb{N}~{}\big{|}~{}\sum_{k\in\mathbb{N}}k\cdot p(k)=m,\sum_{k\in\mathbb{N}}p(k)\leq n\right\}\right|$ (74)
$=\left|\mathsf{Par}(m,\leq_{n})\right|.$ (75)
We are now in a position to prove our lower bounds. Combining (72) and (75), we bound the second moment of the likelihood ratio as follows,
@@
$=\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{n})\right|\rho^{2m}.$ (77)
Since $\left|\mathsf{Par}(m,\leq_{n})\right|$ is a monotone non-decreasing function of $n$ we have
$\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]\leq\sum_{m=0}^{\infty}\left|\mathsf{Par}(m,\leq_{\infty})\right|\rho^{2m},$ (78)
where $|\mathsf{Par}(m,\leq_{\infty})|$ is the number of integer partitions of the number $m$ (with no bound on the number of parts). We comment that bounding $|\mathsf{Par}(m,\leq_{n})|$ by $|\mathsf{Par}(m,\leq_{\infty})|$ (and thereby loosing the decency with $n$) does not compromise our analysis. In Remark Remark we prove a lower bound on $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ which shows that the thresholds obtained cannot be improved by refining the analysis of (77).
In the following, we will use the well-known Hardy-Ramanujan Formula (see, e.g., [29, Chapter 1.3]) for the number of integer partitions, which states that there exists some universal constant $c\in\mathbb{R}$ such that for all $m\in\mathbb{N}\setminus\left\{0\right\}$,
$|\mathsf{Par}(m,\leq_{\infty})|\leq c\cdot\frac{1}{4\sqrt{3}m}\exp\left(\pi\sqrt{\frac{2m}{3}}\right).$ (79)